# Lucas Resende CREST, ENSAE, Institut Polytechnique de Paris, France Guillaume Lecué ESSEC Business School, France Lionel Wilner CREST, ENSAE, Institut Polytechnique de Paris, France Philippe Choné CREST, ENSAE, Institut Polytechnique de Paris, France

### v1

@@ Abstract
Maximum likelihood estimators, such as the Poisson Pseudo-Maximum Likelihood (PPML), suffer from the incidental parameter problem: a bias in the estimation of structural parameters that arises from the joint estimation of structural and nuisance parameters. To address this issue in multi-way gravity models, we propose a novel, asymptotically unbiased estimator. Our method reframes the estimation as a series of classification tasks and is agnostic to both the number and structure of fixed effects. In sparse data environments — common in the network formation literature — it is also computationally faster than PPML. We provide empirical evidence that our estimator yields more accurate point estimates and confidence intervals than PPML and its bias-correction strategies. These improvements hold even under model misspecification and are more pronounced in sparse settings. While PPML remains competitive in dense, low-dimensional data, our approach offers a robust alternative for multi-way models that scales efficiently with sparsity. The method is applied to estimate the effect of a policy reform on spatial accessibility to health care in France.
Keywords weighted networks $\cdot$ incidental parameter problem $\cdot$ gravity models $\cdot$ sparse data $\cdot$ polyad estimator
@@ 1 Introduction
Gravity-type models are extensively used in social sciences to study flows or interactions between entities of different nature. Economists have relied on gravity equations to study bilateral trade flows at least since Timb_1962, see the survey of head2014gravity. Gravity models are relevant in many other fields than trade economics. For instance, they have been applied in labor and development economics to study migration flows between regions or countries, beine2016practitioners; in international finance to model foreign direct investments, see di2005drives; in innovation economics to measure knowledge flows (citations, patents, etc.), peri2005determinants; in health economics to measure interregional patient flows, dingel2023market, as well as spatial accessibility to care, luo2003measures.
@@
Since the important contribution of silva2006log, it is known that the traditional log-linear OLS estimation suffers from bias under heteroskedasticity, particularly when many flows are zero. [^1] Recently, chen2024logs argue that log-like transformations can also distort the interpretation of coefficients as percentage effects, since they depend on the units of the outcome. This observation greatly motivated the adoption of Poisson models for gravity. The properties of Poisson pseudo-maximum likelihood estimators have first been investigated by gourieroux1984pseudo in the absence of fixed effects, with one-way fixed effect panel data applications being pioneered by HHG. To uncover fundamental parameters at disaggregated levels and to control for unobserved heterogeneity along various dimensions, recent applied research has gradually considered models with more levels of fixed effects, introducing higher-dimensional interactions in panel or multi-way gravity models. [^2] Three-way gravity models, with exporter-year, importer-year, and exporter-importer fixed effects, are common in the modern trade literature. Recent studies recognize the economic importance of the sector or product levels, potentially leading to even richer structures of fixed effects, breinlich2024trade and Delb_Dina_WD_2020. As the structure of fixed effects becomes more complex, fernandez2016individual and weidner2021bias show that direct MLE under the Poisson model may be plagued by incidental parameter problems. Specifically, as the sample size grows, so too does the number of nuisance parameters representing the fixed effects, possibly creating bias in standard maximum likelihood estimation of the parameters of interest, as first described by Neym_Scot_Ecma_1948.
@@
In this paper, we tackle the incidental parameter problem in multi-way gravity models by proposing a novel, asymptotically unbiased estimator. In the spirit of graham2017econometric, we regard data through the lens of graph theory. The key insight is that certain configurations of count outcomes within subgroups of observations, which we call polyads, have identical sufficient statistics for the fixed effects, making their relative likelihoods independent of the fixed effects. Our framework differs from graham2017econometric in two important dimensions. First, while Graham models undirected graphs, we consider multipartite graphs. Second, Graham’s network formation model considers only the extensive margin, i.e., the probability that potential links are realized. By contrast, we also consider the intensive margin, i.e., the strength of connections.
@@
Our method compares favorably with existing approaches for gravity models along several dimensions. First, it accommodates multi-way models with an arbitrary number of node groups, in contrast to debiasing methods such as fernandez2016individual and weidner2021bias, which are restricted to two- and three-way structures, and to two-way approaches like graham2017econometric, charbonneau2012multiple, and jochmans2017two. Second, unlike PPML estimation, our estimator is bias-free by construction. Even relative to bias-corrected PPML procedures, our approach remains advantageous: the corrections may themselves be biased in finite samples, as weidner2021bias argue, or computationally prohibitive in large datasets, as in zylkin2024bootstrap. Third, compared with approaches such as charbonneau2012multiple, our estimator better exploits the available variability while remaining computationally feasible, and the convexity of our loss function delivers strong numerical performance relative to more general semiparametric method-of-moments procedures (e.g., jochmans2017two; yang2023three).
@@
Our study is also connected to the literature on discrete choice models for panel and network data (Rasc_1960, Ande_1973,Cham_Hand_1984, Hono_Kyri_Ecma_2000, Magn_Ecma_2004). Presenting the conditional likelihood methods used in these settings, the recent review of Dano_Hono_Weid_WP_2025 highlights how identification strategies relate to difference-in-differences approaches. Specifically, they provide the differencing vectors that are valid to identify the parameters of interest. [^3] Muris_Pakel_WP_2025 follow this approach to study the formation of triadic networks. They introduce a hexad logit estimator that extends the tetrad logit estimator of graham2017econometric. We proceed the same way for count data and gravity models. The polyad estimator can be thought of as a nonlinear version of a difference-in-differences estimator. Contrary to the above cited literature, the polyad method handles count data and recovers both the existence and intensity of relationships. In the special case where the count data is right-censored at one, i.e., where the intensive margin is ignored, the polyad estimator maximizes the same conditional likelihood as graham2017econometric and the studies mentioned in this paragraph .
@@
For researchers working with sparse networks — i.e., where most potential connections are not realized —, our approach offers a distinct computational advantage: polyads can be constructed by looping over pairs of edges with strictly positive counts (i.e., realized connections), allowing the estimation procedure to scale with the number of observed relationships rather than the number of potential relationships. This is a significant step forward, since the usage of tetrad-based methods has been limited by its computational cost: the available methods for computing tetrad-based statistics, which work only for two- or three-way models, either (i) require looping over all pairs of edges, including unrealized connections (e.g., graham2017econometric; Muris_Pakel_WP_2025), or (ii) rely on matrix multiplications that remain slow for sparse networks (e.g., jochmans2017two). In particular, as exemplified by our experiments, our computational implementation enables the use of the polyads method on large administrative datasets. See Section 5.5 for a more detailed discussion.
@@
We establish consistency and asymptotic normality under mild assumptions, extending the current theoretical framework in two main directions: relaxing sampling assumptions and better exploiting convexity. First, we do not require the existence of a distribution to sample the nodes, thus treating the fixed effects as true parameters and revealing fundamental geometrical properties that a graph should satisfy for consistency and asymptotic normality (see Assumptions 2 and 4). Second, we obtain consistency without boundedness assumptions by modifying classical results from newey1994large under the light of convex analysis tools from rockafellar-1970a and asymptotic statistics results from Andersen1982. We also avoid the existence of a limiting risk for asymptotic normality exploring results from victor, which are based on MR1026303; MR1186263.
@@
The practical limitations of our method merit clear statement. First, we do not consider interdependencies between observations beyond those captured by fixed effects and observed covariates. Second, our approach is designed specifically for count data, rather than continuous weights. Third, the method becomes computationally inefficient in dense networks where most potential relationships are realized. Within these constraints, however, our estimator provides a powerful tool for estimating gravity models with high-dimensional fixed effects structures.
@@ 2 Model assumptions
involves a potentially high number of fixed effects. The MLE estimator of the parameter of interest $\beta_{\star}$ has been shown to be asymptotically biased for $D\geq 3$, see weidner2021bias as well as the experiments presented in Section 6.
@@
Taking $D=2$ and $\mathcal{G}=\{\{1\}\}$ yields the structure of the classical model studied by HHG
@@
Taking $D=2$ and $\mathcal{G}=\mathcal{G}^{{\rm max}}=\{\{1\},\{2\}\}$ yields the structure of the standard gravity model studied by silva2006log , i.e., $\ln\lambda_{i_{1}i_{2}}=\beta_{\star}^{\top}X_{i_{1}i_{2}}+\theta_{i_{1}}^{1}+\theta_{i_{2}}^{2}$. The usual econometric model
@@
Taking $D=3$ and $\mathcal{G}=\mathcal{G}^{{\rm max}}=\{\{1,2\},\{1,3\},\{2,3\}\}$ yields the structure of the model studied by weidner2021bias, i.e., $\ln\lambda_{i_{1}i_{2}i_{3}}=\beta_{\star}^{\top}X_{i_{1}i_{2}i_{3}}+\theta_{i_{1}i_{2}}^{1,2}+\theta_{i_{1}i_{3}}^{1,3}+\theta_{i_{2}i_{3}}^{2,3}$. Relabeling the edges $(i_{1},i_{2},i_{3})$ as $(i,j,t)$ and the fixed effects $(\theta_{i_{1}i_{2}}^{1,2},\theta_{i_{1}i_{3}}^{1,3},\theta_{i_{2}i_{3}}^{2,3})$ as $(u_{ij},v_{jt},w_{it})$, we obtain the usual econometric formulation
@@ 3 Estimation of the homophily parameter $\beta_{\star}$ via polyads
To avoid the incidental parameter problem mentioned above, we first condition the likelihood on a sufficient statistics for the fixed effects given by the degrees of the nodes, see Subsection 3.1. This approach has been followed in two-way contexts by charbonneau2012multiple, graham2017econometric and jochmans2018semiparametric. We construct a loss function based on a set of ‘directions’ that generate all the variability in the data for given degrees.
@@ 3.1 Generalized degrees and polyad transformations
which we call the degree of $\rho$ relative to the fixed effect level $g$ in the graph $y$. This quantity generalizes the notion of degree used in graham2017econometric, capturing the total number of connections between edges $\bf i$} for which $g({\bf i})=\rho$. [^4] Consider for instance Example 3, where the index $i_{1},i_{2},i_{3}$ are denoted $i,j,t$ as in many gravity models. The degree of (4,5) relative to the fixed effect level $u_{ij}$ is $\delta^{1,2}_{4,5}(y)=\sum_{t}y_{45t}$. The set of degrees for the fixed effects level $g\in\mathcal{G}$ is denoted by $\delta^{g}(y)=\left[\delta^{g}_{\rho}(y)\right]_{\rho\in g(\mathcal{I})}$. The set of all degrees is denoted by $\delta(y)=\left[\delta^{g}(y)\right]_{g\in\mathcal{G}}$.
@@
Polyads as introduced in Definition 1 are generalizations to the $D$-dimensional framework of tetrads from charbonneau2012multiple; graham2017econometric; jochmans2018semiparametric defined for $D=2$. The total number of polyads, i.e. the size of $\Xi$, is $\prod_{d=1}^{D}n_{d}(n_{d}-1)$. In the upper left corner of Figure 1 we exemplify a polyad on $D=2$,
@@ 3.2 A classification problem on polyads and the associated estimator of $\beta_{\star}$
The above polyad estimator can be interpreted in relation to a logit classification problem. Specifically, given a polyad $\xi$ and the observed orbit $\mathcal{O}_{\xi}(Y)$, the minimal rank $m_{\xi}(Y)$ is distributed according to a conditional Logit model, see McFa_Chap_1974. To see this, consider the transformed graph $\underline{y}=T^{-m_{\xi}(Y)}(Y)$, which can be thought of as the “minimal” graph in the orbit of the observed graph $y$. In the example of Figure 1, the graph $\underline{y}$ is the second graph among the six graphs shown on the bottom line (that is for $r=-m_{\xi}(y)=-1$). The observed orbit $\mathcal{O}_{\xi}(Y)$ can thus be represented as
@@
(Analogy with conditional likelihood methods used in panel and network data) For any active polyads $\xi\in\Xi_{a}$, the tensor $(s_{\xi}({\bf i}))_{\bf i}$, whose entries belong to $\{-1,0,1\}$, is a “differencing vector” in the sense of Dano_Hono_Weid_WP_2025. It applies linearly to the independent features $X$ to deliver the generalized difference-in-differences term $\widetilde{X}_{\xi}$. The corresponding transformation $T_{\xi}$ plays the same role for the count data $Y$. Conditionally on the observed graph $y$ belonging to the orbit $\mathcal{O}_{\xi}(Y)$, we thus obtain a conditional logit classification problem, with the number of alternatives, $|\mathcal{O}_{\xi}(Y)|-1$, being polyad-specific.
@@ 4 Large sample properties of the polyads estimator
Recall that $n=\prod_{d=1}^{D}n_{d}$ is the size of the graph indexed by $\mathcal{I}=[n_{1}]\times[n_{2}]\times\dots\times[n_{D}]$ and that if $\Xi$ is the set of polyads given by the indices $\mathcal{I}$, then $|\Xi|$ has order $n^{2}$. Through this section, one is given a sequence $\left(\Xi^{(1)},\Xi^{(2)},\dots\right)$ of families of polyads, defined over growing sets of indices. Let $n_{d}^{(k)}$ be the size of the $d$-th dimension associated with $\Xi^{(k)}$. We impose no restrictions on how each dimension $n_{d}^{(k)}$ grows, requiring only that $n^{(k)}=\prod_{d=1}^{D}n_{d}^{(k)}\to\infty$ — or, equivalently, $|\Xi^{(k)}|\to\infty$. Notably, our results accommodate short panels where one dimension remains bounded while others diverge. For three-way models, this includes settings where the PPML estimator suffers from the incidental parameter problem (weidner2021bias). We also do not assume the existence of a limiting function $Q_{\infty}$, thus avoiding restrictive assumptions on the asymptotic behavior of fixed effects and covariates. In particular, we do not impose that fixed effects or covariates are drawn from any probability distribution. To circumvent a convoluted notation we use $(k)$ as index in place of $\Xi^{(k)}$, for instance, $\widehat{\beta}_{\Xi^{(k)}}$ is written as $\widehat{\beta}^{(k)}$. Also, notice that associated with each $(k)$ we also have covariates $X^{(k)}$, fixed effects $\theta^{(k)}$ and random variables $Y^{(k)}$.
Finally, unlike graham2017econometric; jochmans2018semiparametric, we do not require the components of $\left[n_{d}^{(k)}\right]$ to arise from random sampling. This permits a more general interpretation of the data generating process. Consider a two-way model of doctor-patient interactions. Under the framework of graham2017econometric; jochmans2018semiparametric, one assumes the existence of a large population graph containing all doctors and patients, from which patients and doctors are randomly sampled, inducing distributions on both fixed effects and covariates. Our approach is more constructive: for each $(k)$, the distribution of fixed effects and covariates may differ entirely. In the doctor-patient example, our framework accommodates data collection that expands geographically. For instance, initially observing patients and doctors in Paris only, then adding doctors from Marseilles, then adding also patients from Marseilles, and so forth. Imposing a random sampling structure would limit this possibility, as any finite sample could contain doctors from Marseilles with positive probability.

### v2

# Statistical inference in large multi-way networks

@@ Abstract
We propose a new method to estimate structural parameters in multi-way networks while controlling for rich structures of fixed effects. The method is based on a series of classification tasks and is agnostic to both the number and structure of fixed effects. In contrast to full maximum likelihood approaches, our estimator does not suffer from the incidental parameter problem. For sparsely connected networks, it is also computationally faster than PPML. We provide empirical evidence that our estimator yields more reliable confidence intervals than PPML and its bias-correction strategies. These improvements hold even under model misspecification and are more pronounced in sparse settings. While PPML remains competitive in dense, low-dimensional data, our approach offers a robust alternative for multi-way models that scales efficiently with sparsity. The method is applied to study the causal effect of a policy reform on spatial accessibility to health care in France.
Keywords multi-way networks $\cdot$ polyadic data $\cdot$ high-dimensional fixed effects $\cdot$ incidental parameter problem $\cdot$ gravity models $\cdot$ sparse networks $\cdot$ weighted networks $\cdot$ conditional likelihood
@@ 1 Introduction
Network data, which are becoming available at increasingly granular levels, are receiving a great deal of attention in economics, see Graham and De Paula (2020). In this paper, we consider “multi-way” networks that involve interactions between entities of different nature such as importing and exporting countries, buyers and suppliers, teachers and schools, doctors and patients. The strength of the interactions in such networks is commonly measured at disaggregated levels, e.g., industries or products for trade data, consultations and medical procedures for health data, patents and citations for innovation data, etc. Multi-way network data, sometimes referred to as “polyadic”, are indexed by multidimensional indices that represent the relevant dimensions in each case, for instance exporter, importer, product, and time in the trade example.
@@
To model connections in multi-way networks and control for unobserved heterogeneity along various dimensions, recent applied research has gradually considered models with richer structures of fixed effects, involving higher-dimensional interactions. For instance, three-way gravity models, with exporter-year, importer-year, and exporter-importer fixed effects, are common in the modern trade literature. [^1] Recent studies recognize the economic importance of the sector or product levels, potentially leading to even richer structures of fixed effects, Breinlich et al. (2024) and Delbourg and Dinar (2020). Yet as the structure of fixed effects becomes more complex, maximum likelihood estimators may be plagued by incidental parameter problems, see Fernández-Val and Weidner (2016) and Weidner and Zylkin (2021). Specifically, as the sample size grows, so too does the number of nuisance parameters representing the fixed effects, possibly creating bias in standard maximum likelihood estimation of the parameters of interest, as first described by Neyman and Scott (1948).
@@
In this paper, we propose a novel estimator that does not suffer from the incidental parameter problem and handles multi-way gravity models. In the spirit of Graham (2017), we regard data through the lens of graph theory. The key insight is that certain configurations of outcomes within subgroups of observations, which we call polyads, have identical sufficient statistics for the fixed effects, making their relative likelihood independent of the fixed effects. Our framework differs from Graham (2017) in two important dimensions. First, while Graham models undirected graphs, we use multipartite graphs to model multi-way networks. Second, Graham’s network formation model considers only the extensive margin, i.e., the probability that potential links are realized. By contrast, we study the strength of connections in weighted networks, thus modeling both the intensive and extensive margins.
@@
This study is connected to the strand of the gravity literature starting with Silva and Tenreyro (2006). [^2] Their seminal paper shows that traditional log-linear OLS estimation suffers from bias under heteroskedasticity, particularly when many flows are zero, which greatly motivated the adoption of Poisson models for gravity. The properties of Poisson pseudo-maximum likelihood estimators have first been investigated by Gourieroux et al. (1984) in the absence of fixed effects, with one-way fixed effect panel data applications being pioneered by Hausman et al. (1984). Recently, Chen and Roth (2024) argue that log-like transformations can also distort the interpretation of coefficients as percentage effects, since they depend on the units of the outcome. Our method compares favorably with recent econometric studies along several dimensions. First, it accommodates multi-way models with an arbitrary number of node groups, in contrast to debiasing methods such as Fernández-Val and Weidner (2016), Jochmans (2017), and Weidner and Zylkin (2021), which are restricted to two- and three-way structures. Second, unlike PPML estimation, our estimator has no incidental parameter problem by construction. Even relative to bias-corrected PPML procedures, our approach remains advantageous: the corrections may themselves be biased in finite samples, as Weidner and Zylkin (2021) argue, or computationally prohibitive in large datasets, as in Zylkin (2024). Third, compared with approaches such as Charbonneau (2012), our estimator better exploits the available variability while remaining computationally feasible, and the convexity of our loss function delivers strong numerical performance relative to more general semiparametric method-of-moments procedures (e.g., Jochmans (2017); Yang and Zhang (2023)).
@@
Our study is also connected to the literature on discrete choice models for panel and network data (Rasch (1960), Ande_JRSS_1970,Chamberlain (1984), Honoré and Kyriazidou (2000), Magnac (2004)). Presenting the conditional likelihood methods used in these settings, the recent review of Dano et al. (2025) highlights how identification strategies relate to difference-in-differences approaches. Specifically, they provide the differencing vectors that are valid to identify the parameters of interest. [^3] Muris and Pakel (2025) follow this approach to study the formation of triadic networks. They introduce a hexad logit estimator that extends the tetrad logit estimator of Graham (2017). We proceed the same way for count data and multi-way networks. The polyad estimator can be thought of as a nonlinear version of a difference-in-differences estimator. Contrary to the above cited literature, the polyad method handles count data and recovers both the existence and intensity of relationships. In the special case where the count data is right-censored at one, i.e., where the intensive margin is ignored, the polyad estimator maximizes the same conditional likelihood as Graham (2017) and the studies mentioned in this paragraph .
@@
For researchers working with sparse networks — i.e., where most potential connections are not realized —, our approach offers a distinct computational advantage: polyads can be constructed by looping over pairs of edges with strictly positive counts (i.e., realized connections), allowing the estimation procedure to scale with the number of observed relationships rather than the number of potential relationships. This is a significant step forward, since the usage of tetrad-based methods has been limited by its computational cost: the available methods for computing tetrad-based statistics, which work only for two- or three-way models, either (i) require looping over all pairs of edges, including unrealized connections (e.g., Graham (2017); Muris and Pakel (2025)), or (ii) rely on matrix multiplications that remain slow for sparse networks (e.g., Jochmans (2017)). In particular, as exemplified by our experiments, our computational implementation enables the use of the polyads method on large administrative datasets. See Section 5.5 for a more detailed discussion.
@@
We establish consistency and asymptotic normality under mild assumptions, extending the current theoretical framework in two main directions: relaxing sampling assumptions and better exploiting convexity. First, unlike Graham (2017, 2020), we do not require the existence of a distribution to sample the nodes, thus treating the fixed effects as true parameters and revealing fundamental geometrical properties that a graph should satisfy for consistency and asymptotic normality (see Assumptions 2 and 4). Second, we obtain consistency without boundedness assumptions by modifying classical results from Newey and McFadden (1994) under the light of convex analysis tools from Rockafellar (1970) and asymptotic statistics results from Andersen and Gill (1982). We also avoid the existence of a limiting risk for asymptotic normality exploring results from Brunel (2020), which are based on Haberman (1989); Niemiro (1992).
@@
The practical limitations of our method merit clear statement. First, we do not consider interdependencies between observations beyond those captured by fixed effects and observed covariates. Second, our approach is designed specifically for count data, rather than continuous weights. Third, the method becomes computationally inefficient in dense networks where most potential relationships are realized. Within these constraints, however, our estimator provides a powerful tool for inference in multi-way networks with high-dimensional fixed effects structures.
@@ 2 Model assumptions
involves a potentially high number of fixed effects. The MLE estimator of the parameter of interest $\beta_{\star}$ has been shown to be asymptotically biased for $D\geq 3$, see Weidner and Zylkin (2021) as well as the experiments presented in Section 6.
@@
Taking $D=2$ and $\mathcal{G}=\{\{1\}\}$ yields the structure of the classical model studied by Hausman et al. (1984)
@@
Taking $D=2$ and $\mathcal{G}=\mathcal{G}^{{\rm max}}=\{\{1\},\{2\}\}$ yields the structure of the standard gravity model studied by Silva and Tenreyro (2006) , i.e., $\ln\lambda_{i_{1}i_{2}}=\beta_{\star}^{\top}X_{i_{1}i_{2}}+\theta_{i_{1}}^{1}+\theta_{i_{2}}^{2}$. The usual econometric model
@@
Taking $D=3$ and $\mathcal{G}=\mathcal{G}^{{\rm max}}=\{\{1,2\},\{1,3\},\{2,3\}\}$ yields the structure of the model studied by Weidner and Zylkin (2021), i.e., $\ln\lambda_{i_{1}i_{2}i_{3}}=\beta_{\star}^{\top}X_{i_{1}i_{2}i_{3}}+\theta_{i_{1}i_{2}}^{1,2}+\theta_{i_{1}i_{3}}^{1,3}+\theta_{i_{2}i_{3}}^{2,3}$. Relabeling the edges $(i_{1},i_{2},i_{3})$ as $(i,j,t)$ and the fixed effects $(\theta_{i_{1}i_{2}}^{1,2},\theta_{i_{1}i_{3}}^{1,3},\theta_{i_{2}i_{3}}^{2,3})$ as $(u_{ij},v_{jt},w_{it})$, we obtain the usual econometric formulation
@@ 3 Estimation of the homophily parameter $\beta_{\star}$ via polyads
To avoid the incidental parameter problem mentioned above, we first condition the likelihood on a sufficient statistics for the fixed effects given by the degrees of the nodes, see Subsection 3.1. This approach has been followed in two-way contexts by Charbonneau (2012), Graham (2017) and Jochmans (2018). We construct a loss function based on a set of ‘directions’ that generate all the variability in the data for given degrees.
@@ 3.1 Generalized degrees and polyad transformations
which we call the degree of $\rho$ relative to the fixed effect level $g$ in the graph $y$. This quantity generalizes the notion of degree used in Graham (2017), capturing the total number of connections between edges $\bf i$} for which $g({\bf i})=\rho$. [^4] Consider for instance Example 3, where the index $i_{1},i_{2},i_{3}$ are denoted $i,j,t$ as in many gravity models. The degree of (4,5) relative to the fixed effect level $u_{ij}$ is $\delta^{1,2}_{4,5}(y)=\sum_{t}y_{45t}$. The set of degrees for the fixed effects level $g\in\mathcal{G}$ is denoted by $\delta^{g}(y)=\left[\delta^{g}_{\rho}(y)\right]_{\rho\in g(\mathcal{I})}$. The set of all degrees is denoted by $\delta(y)=\left[\delta^{g}(y)\right]_{g\in\mathcal{G}}$.
@@
Polyads as introduced in Definition 1 are generalizations to the $D$-dimensional framework of tetrads from Charbonneau (2012); Graham (2017); Jochmans (2018) defined for $D=2$. The total number of polyads, i.e. the size of $\Xi$, is $\prod_{d=1}^{D}n_{d}(n_{d}-1)$. In the upper left corner of Figure 1 we exemplify a polyad on $D=2$,
@@ 3.2 A classification problem on polyads and the associated estimator of $\beta_{\star}$
The above polyad estimator can be interpreted in relation to a logit classification problem. Specifically, given a polyad $\xi$ and the observed orbit $\mathcal{O}_{\xi}(Y)$, the minimal rank $m_{\xi}(Y)$ is distributed according to a conditional Logit model, see McFadden (1974). To see this, consider the transformed graph $\underline{y}=T^{-m_{\xi}(Y)}(Y)$, which can be thought of as the “minimal” graph in the orbit of the observed graph $y$. In the example of Figure 1, the graph $\underline{y}$ is the second graph among the six graphs shown on the bottom line (that is for $r=-m_{\xi}(y)=-1$). The observed orbit $\mathcal{O}_{\xi}(Y)$ can thus be represented as
@@
(Analogy with conditional likelihood methods used in panel and network data) For any active polyads $\xi\in\Xi_{a}$, the tensor $(s_{\xi}({\bf i}))_{\bf i}$, whose entries belong to $\{-1,0,1\}$, is a “differencing vector” in the sense of Dano et al. (2025). It applies linearly to the independent features $X$ to deliver the generalized difference-in-differences term $\widetilde{X}_{\xi}$. The corresponding transformation $T_{\xi}$ plays the same role for the count data $Y$. Conditionally on the observed graph $y$ belonging to the orbit $\mathcal{O}_{\xi}(Y)$, we thus obtain a conditional logit classification problem, with the number of alternatives, $|\mathcal{O}_{\xi}(Y)|-1$, being polyad-specific.
@@ 4 Large sample properties of the polyads estimator
Recall that $n=\prod_{d=1}^{D}n_{d}$ is the size of the graph indexed by $\mathcal{I}=[n_{1}]\times[n_{2}]\times\dots\times[n_{D}]$ and that if $\Xi$ is the set of polyads given by the indices $\mathcal{I}$, then $|\Xi|$ has order $n^{2}$. Through this section, one is given a sequence $\left(\Xi^{(1)},\Xi^{(2)},\dots\right)$ of families of polyads, defined over growing sets of indices. Let $n_{d}^{(k)}$ be the size of the $d$-th dimension associated with $\Xi^{(k)}$. We impose no restrictions on how each dimension $n_{d}^{(k)}$ grows, requiring only that $n^{(k)}=\prod_{d=1}^{D}n_{d}^{(k)}\to\infty$ — or, equivalently, $|\Xi^{(k)}|\to\infty$. Notably, our results accommodate short panels where one dimension remains bounded while others diverge. For three-way models, this includes settings where the PPML estimator suffers from the incidental parameter problem (Weidner and Zylkin, 2021). We also do not assume the existence of a limiting function $Q_{\infty}$, thus avoiding restrictive assumptions on the asymptotic behavior of fixed effects and covariates. In particular, we do not impose that fixed effects or covariates are drawn from any probability distribution. To circumvent a convoluted notation we use $(k)$ as index in place of $\Xi^{(k)}$, for instance, $\widehat{\beta}_{\Xi^{(k)}}$ is written as $\widehat{\beta}^{(k)}$. Also, notice that associated with each $(k)$ we also have covariates $X^{(k)}$, fixed effects $\theta^{(k)}$ and random variables $Y^{(k)}$.
Finally, unlike Graham (2017); Jochmans (2018), we do not require the components of $\left[n_{d}^{(k)}\right]$ to arise from random sampling. This permits a more general interpretation of the data generating process. Consider a two-way model of doctor-patient interactions. Under the framework of Graham (2017); Jochmans (2018), one assumes the existence of a large population graph containing all doctors and patients, from which patients and doctors are randomly sampled, inducing distributions on both fixed effects and covariates. Our approach is more constructive: for each $(k)$, the distribution of fixed effects and covariates may differ entirely. In the doctor-patient example, our framework accommodates data collection that expands geographically. For instance, initially observing patients and doctors in Paris only, then adding doctors from Marseilles, then adding also patients from Marseilles, and so forth. Imposing a random sampling structure would limit this possibility, as any finite sample could contain doctors from Marseilles with positive probability.